OpenStudy (anonymous):

Move everything away from y by using the additive inverse/opposite 5x + 2y = 8

OpenStudy (anonymous):

5x+2y=8 2y=-5x+8 y=-5x/2+4

OpenStudy (muhammad_nauman_umair):

5x-5x+2y=8-5x 2y/2=(8-5x)/2 y=4-5/2x